OpenStudy (anonymous):

Tan (x/2)= square root of 3 so far I have x can = π/3 and 4π/7 I am lost to do next.

OpenStudy (anonymous):

4π/3 I mean

OpenStudy (loser66):

Your original is (x/2) not x, hence for (x/2) = pi/3, x = 2pi/3 for (x/2) = 4pi/3, x = 8pi/3 2pi/3 and 8pi/3 has the same destination.